OpenStudy (hellomynameiskam):

A bag contains 12 red marbles and 4 blue marbles. What is the ratio, written in lowest terms, of red marbles to the total number of marbles in the bag? A. 1 to 4 B. 1 to 3 C. 2 to 3 D. 3 to 4

HanAkoSolo (jamierox4ev3r):

Hi there! So there are 12 red marbles in the bag, then how many total marbles are there? Remember, ratios are simply fractions

OpenStudy (anonymous):

how many marbles are in the bag?

OpenStudy (hellomynameiskam):

16 marbles are in the bag

HanAkoSolo (jamierox4ev3r):

right. So therefore, if you're trying to find a ratio of red marbles to total marbles, you would have a numerator of the number of red marbles (12) and a denominator of total amount of marbles (16)

OpenStudy (anonymous):



HanAkoSolo (jamierox4ev3r):

So thus far, your ratio would look like this: \(\Large\frac{12}{16}\)

OpenStudy (anonymous):

ur welcome

HanAkoSolo (jamierox4ev3r):

@qwertyboy56 Please DO NOT provide direct answers. It is against the rules here

OpenStudy (anonymous):

oh srry

HanAkoSolo (jamierox4ev3r):

@HELLOMYNAMEISKAM you following me so far? Once you have the fraction of \(\Large\frac{12}{16}\) , then this can be simplified to find the answer

OpenStudy (anonymous):

its right tho

OpenStudy (hellomynameiskam):

yeah. @Jamierox4ev3r

OpenStudy (anonymous):

true... srry

HanAkoSolo (jamierox4ev3r):

Don't feel too bad, just as long as you don't do it again :) and as for the problem, the fastest way to simplify this is to find a greatest common factor. What is a number that you could divide both 12 and 16 by?

OpenStudy (hellomynameiskam):

2

HanAkoSolo (jamierox4ev3r):

Factors of 12: 1,2,3,4,12 Factors of 16: 1,2,4,8,16 Yes, 2 is a factor that both of these numbers share. However, 4 is a factor that is greater and works for more numbers, so this will probably lead us to the correct answer faster Does that make sense?

OpenStudy (hellomynameiskam):

yh

HanAkoSolo (jamierox4ev3r):

alright :) so if you divide 12 by 4, and 16 by 4, which fraction do you get?

OpenStudy (hellomynameiskam):

i got 12/4=3 16\4=4

HanAkoSolo (jamierox4ev3r):

yep. so that would give you a fraction of \(\Large\frac{3}{4}\). And that is the final simplified ratio :)

OpenStudy (hellomynameiskam):

ok thanks